Whats an example of a system of equation that is more easily solved by substitution, and a system of equations that is more easily solved by elimination?
Need 2 equations please! :)

Answer by fractalier(6550)

y = 3x + 1
5x - 2y = 8
is easier solved by substituting in for y in the second equation, whereas
3x + 4y = 11 and
6x - 4y = 16
is easier solved by the elimination method...when you add them, the 4y terms drop out...